A garden plot is to contain 128 sq. ft. If its length is to be twice its width, what should its dimensions be?



1
Expert's answer
2021-01-04T20:11:39-0500

It is given that garden plot is to contain "128" sq.ft .

Therefore area of the garden plot is "128" sq.ft.

According to the problem length of the garden plot is twice its width.

Let width of the garden plot is "l" ft.

Then length of the garden plot will be "2l" ft.

Therefore area of the garden plot is

"= 2l\u00d7l" sq.ft.

"=2l^2" sq.ft.

According to the problem,

"2l^2=128"

"\implies l^2=64"

"\implies l=\sqrt{64}"

"\implies l=8"

So width of the garden plot is "= 8" ft.

Therefore length of the garden plot is "=16" ft.

Therefore dimensions of the garden plot is "16" ft. by "8" ft.